OpenStudy (anonymous):

How can you graph conics (parabola, ellipse, hyperbola)? I know they have formulas... I just forgot.

OpenStudy (phi):

That is a pretty open question. Do you mean by hand? Or by using a graphing calculator? Or using a computer (try wolfram (google it))?

OpenStudy (anonymous):

manually. by hand. We're not allowed to use calculators in exams anyway

OpenStudy (phi):

In that case, learn what the equation looks like for each type. for example, an ellipse is \[ \frac{x^2}{a^2} + \frac{y^2}{b^2} =1 \] It is centered at 0,0. It looks like an oval. pick some easy points. if x=0, then y= ±b (you need to be able to figure that out) . that is (0,+b), and (0,-b) plot those points. Do the same for y=0 . you get (+a,0) and (-a,0). connect the dots with a smooth curve. Use wolfram to check your answers, or to see what it should look like.

OpenStudy (anonymous):

Yeah this is pretty open question. I can help with ellipses. Before drawing ellipses you need to know how to draw circles. The most simple drawing of a circle is X^2 + Y^2 = 1 That's the unit circle. From there you add coefficients to x and y in order to strech the circle out isentially drawing an ellipse.

OpenStudy (anonymous):

Thanks I'll try the doing the ellipse